OpenStudy (anonymous):

(x + 1) / (x^2 -1) + (x-1) / (x + 1) = 1 / (x - 1) can you tell me the value of x an show me how you found it please

OpenStudy (anonymous):

the expression is equal to 1/(x-1) + (x-1)/(x+1) =1/(x-1) the right hand term cancels out leaving only the centre term provided x is not -1 we can say that x-1 = x+ 1 => -1 = 1 which is clearly not true ...therfore x must be 0

jhonyy9 (jhonyy9):

the first step in what you need to do one substitution that x not can being equal with +/- 1 because than this fraction has denominators zero and hence this fractions are invalids